# Wormhole Formation with Gravitational Particle Creation Mechanism

### v2

@@ II Wormhole Solution from Particle Creation Mechanism
Figure 1: The flareout and asymptotic conditions are shown for a few combinations of the model parameter $\beta$ and the proportionality constant $K$. Here throat radius $r_{0}$ is assumed to be 1.
The shape function given by Eq. (13) determines the configuration of the wormhole which is sustained by the negative particle creation pressure. Interestingly, the wormhole does not need an ad-hoc introduction of exotic fluids to account for the negative pressure required to sustain the wormhole geometry. To explore the physical nature of the matter content inside the wormhole geometry, we shall make use of the energy conditions. A generic aspect of wormhole systems is the violation of the null energy condition (NEC) (i.e. $\rho+p_{r}\geq 0$ or $\rho+p_{t}(r)\geq 0$ in terms of radial and tangential pressures respectively) at the throat of the wormhole structure. On the other hand, for some specific models keeping valid the NEC while keeping the dominant energy condition (DEC) ($\rho-p_{r}(r)\geq 0$ or $\rho-p_{t}(r)\geq 0$) violated at the throat is also possible [35, 36, 37, 38, 39]. In this work, we shall construct traversable wormhole structures with the violation of the NEC as well as DEC at the throat of the wormhole, simultaneously keeping the energy condition $\rho\geq 0$. This turns out, as a result of the violation of the weak energy condition (WEC) ($\rho\geq 0,\rho+p_{r}(r)\geq 0\text{ or }\rho+p_{t}(r)\geq 0$). From Eq. (10), given the shape function (13), one may obtain the energy density, radial and tangential pressures given by
$\rho(r)=\frac{K}{\beta r^{2}(\omega+1)},$ (14)
$p_{r}(r)=\frac{K\left(r-\frac{8\pi(r-r_{0})}{\beta(\omega+1)}\right)-r_{0}}{r^{3}},$ (15)
$p_{t}(r)=\frac{2Kr-\frac{8\pi Kr_{0}}{\beta\omega+\beta}+r_{0}}{2r^{3}},$ (16)
To study the energy conditions for the wormhole geometry (13), let us calculate the quantities $\rho+p_{r}(r)$, $\rho+p_{t}(r)$, $\rho-p_{r}(r)$ and $\rho-p_{t}(r)$ given by
$\rho+p_{r}(r):=\frac{Kr(\beta\omega+\beta-8\pi+1)+8\pi Kr_{0}-\beta r_{0}(\omega+1)}{\beta r^{3}(\omega+1)},$ (17)
$\rho+p_{t}(r):=\frac{2K\left(\frac{r-4\pi r_{0}}{\beta\omega+\beta}+r\right)+r_{0}}{2r^{3}},$ (18)
and
$\rho-p_{r}(r):=\frac{\frac{K(8\pi r+r-8\pi r_{0})}{\beta(\omega+1)}-Kr+r_{0}}{r^{3}}.$ (19)
$\rho-p_{t}(r):=-\frac{2K\left(r-\frac{r+4\pi r_{0}}{\beta\omega+\beta}\right)+r_{0}}{2r^{3}}.$ (20)
Figure 2: The NEC and DEC for the (13) is shown for three values of $K$ and $\beta$
Fig. 2 displays the behaviour of the NEC and DEC for different combinations of the model parameters. The NEC in terms of radial pressure shows clear violation at the throat but obeys at far away regions from the throat, whereas the DEC is obeyed at the throat as well as at far away regions. However, in terms of the tangential pressure, the NEC is obeyed at the throat as well as in asymptotic regions but DEC is violated at the throat and satisfied at the asymptotic regions. Also as $\rho\geq 0$, one can say WEC is violated for the radial pressure case and obeyed for the tangential pressure one.
III Toy Shape Functions and Evolution of Particle Creation Pressure
In the previous section, we utilized a phenomenological variation law of the particle creation pressure and found a shape function corresponding to that scenario. In this section, let us assume some specific shape functions and study the evolution of the particle creation pressure.
III.1 Shape function 1: $b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$
@@
which clearly is a positive function for all values of $r$. From Fig.3, it is seen that the shape function obeys the viability conditions mentioned in section II for different values of $\epsilon$.
Refer to caption Refer to caption
Figure 3: The flaring out and asymptotic flatness condition for the shape function (21) is shown for different combinations of the model parameters.
@@
Also, the quantities related to the energy conditions are found to be
@@
Figure 4: NEC and DEC for the shape function (21) for different combination of the model parameters
From Fig. 4, it is observed that the NEC is violated in terms of radial pressure shown at the throat but tends to obey at the asymptotic limit, whereas the DEC is obeyed at the throat as well as at the asymptotic limit. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is obeyed at the throat for the parameter values $(\beta,\epsilon)=(0.2,0.10)\text{ and }(0.5,0.15)$, however, crosses into the violation region at around $r=3$ and $r=1.6$ respectively for the ranges we considered. For $(\beta,\epsilon)=(0.7,0.20)$, the DEC is always violated. Also, as $\rho\geq 0$, WEC is violated for the radial pressure case and obeyed for the tangential pressure one. Moreover, from Eq. (25) and Fig. 5, it is evident that the particle creation pressure is negative at the throat all the way to $r\to\infty$, which is required and expected for backreaction pressure responsible for sustaining a wormhole geometry.
@@ III.2 Shape function 2: $b(r)=r_{0}^{n}r^{1-n}$
where $n$ is a model parameter and $r_{0}$ is the throat radius. This type of shape function has been studied in [41], in the context of $f(R)$ gravity reconstruction for given choices of wormhole shape functions. Fig. 6 clearly shows the required criteria satisfied for viability for different combinations of the associated model parameters.
Refer to captionRefer to caption
Figure 6: The flaring out and asymptotic flatness condition for the shape function (30) is shown for different combinations of the model parameters.
@@
Figure 7: NEC and DEC for the shape function (30)
@@
Fig. 7, shows that the NEC is violated in terms of radial pressure at the throat but tends to obey as $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is violated at the throat and obeyed at $r\to\infty$. Also, WEC is violated in terms of the radial pressure but obeyed in terms of the tangential pressure. Moreover, from Eq. (34) and Fig. 8 (left), it is seen that the particle creation pressure is negative at the throat to $r\to\infty$. Similar observation is drawn for the shape function (39), given by Eq. (43) and shown in Fig. 8 (right) the particle creation pressure tends to zero as $r\to\infty$ like all the other cases.
III.3 Shape function 3: $b(r)=r\exp[\Omega(r-r_{0})]$
@@
where $\Omega$ is the only model parameter and $r_{0}$ being the throat radius. This type of shape function has been considered widely in Ref. [42, 43, 44] in both GR and modified gravity. Fig. 9 shows the validity of the flaring-out and asymptotic flatness for different combinations of the associated model parameters.
Refer to captionRefer to caption
Figure 9: The flaring out and asymptotic flatness condition for the shape function (39) is shown for different combinations of the model parameters.
The energy density corresponding to this shape function is
@@
$p_{r}(r)=\frac{\beta(\omega+1)\left(e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}\right)}{8\pi r^{2}}-\frac{e^{-(\Omega(r-r_{0}))}}{r^{2}},$ (41)
$p_{t}(r)=\frac{e^{\Omega(r_{0}-r)}(4\pi r\Omega-\beta(\omega+1)(r\Omega-1))}{8\pi r^{2}},$ (42)
@@
$\Pi(r)=-\frac{\beta(\omega+1)\left(e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}\right)}{8\pi r^{2}}.$ (43)
@@
$\rho+p_{r}(r):=-\frac{e^{\Omega(r_{0}-r)}((\beta\omega+\beta+1)(r\Omega-1)+8\pi)}{8\pi r^{2}},$ (44)
@@
Figure 10: NEC and DEC for the shape function (39) for different combinations of model parameters.
From Fig. 10, we see NEC violation in terms of radial pressure at the throat but is obeyed at $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is satisfied at the throat as well as at asymptotic regions but DEC remains violated at the throat but obeyed at $r\to\infty$. Moreover, WEC is violated for the radial pressure case and satisfied for the tangential pressure one. Also, from Eq. (43) and Fig. 8 (right), it is seen that the particle creation pressure is negative at the throat and tends to zero at $r\to\infty$.
IV 3D Embedding
$ds^{2}=\left(1-\frac{b(r)}{r}\right)^{-1}dr^{2}+r^{2}d\phi^{2}.$ (48)
The metric (48) is equivalent to
$ds^{2}=dz^{2}+dr^{2}+r^{2}d\phi^{2},$ (49)
which can now be compared with Eq. (48) to obtain
$\frac{dz}{dr}=\pm\left(\frac{r}{b(r)}-1\right)^{-1/2},$ (51)
Eq. (51) can be solved to obtain a functional form of $z(r)$. The shape function obtained from the physical requirements of particle creation given by (13) is shown in the Fig. 11 for $\beta=0.7$ and $K=0.016$.
Figure 11: The 3D embedding of the wormholes are shown for the shape functions (13) the throat is assumed at $r_{0}=1$ and the values of $\beta$ and $K$ is taken to be $0.7$ and $0.016$ respectively.
V Conclusion
@@
The concept of gravitational particle creation phenomena of open thermodynamic cosmological systems, was proposed as a potential explanation to the late-time accelerated expansion of the Universe and inflationary expansion during the early stages of the Universe. This mechanism involves the generation of repulsive gravitational effects, resulting in negative pressure, which could explain the accelerated expansion of the Universe in its later stages. Gravitational particle creation allows for speculation about violations of the NEC, which is crucial for the maintenance of a wormhole structure. This study explores the hypothesis that the geometry of a wormhole can be supported by the negative pressure created by particles. The initial section of the paper focuses on deriving the shape function of a wormhole based on a specific form of particle creation pressure. Specifically, we proposed an inverse-squared variation of particle creation pressure with the radial coordinate, with a source function that is linear in the Hubble parameter at the current time, denoted by $\Gamma=3\beta H_{0}$ with $\beta$ as a free parameter. Through out the paper, owing to the finiteness of the redshift function, we considered a constant redshift function for the analysis. The shape function obtained seems to adhere to the required criteria for a viable wormhole structure and violates the NEC with a positive energy density. In the later part of the paper, in the context of particle creation mechanism we obtained the radial variation of energy density, radial and tangential pressure, and the particle creation pressure within the wormhole geometry, by assuming a few well known shape functions. The radial NEC is found to be violated at the throat in all the cases, which resonates with the behaviour of so-called exotic matter but without the need of introducing explicit form of exotic matter density like chaplygin gas [45, 39, 46, 47, 48, 49, 50, 51] or phantom matter [52, 53, 54, 55, 56, 57, 58, 59, 60]. Finally, the 3D embedding for the wormhole solution obtained from the particle creation phenomena is also carried out.
@@
This study explores the possibility for developing a viable wormhole geometry supported by the back reaction pressure generated from the irreversible particle creation, within an open thermodynamic system. Nevertheless, our work does not address the effect of particle creation on particle geodesics within the wormhole structure or related aspects, as it is beyond the current study’s scope. This prospect will be further investigated in the future work.

### v3

@@ IV Embedding
Figure 3: The 2D and 3D embedding of the wormholes are shown for the shape functions (20) the throat is assumed at $r_{0}=1$ and the values of $\beta$ and $K$ is taken to be $8.55$ and $0.005$ respectively.
The embedding function and the diagrams for the toy models (37),(47) and (56) can be found in [72, 73, 74].
V Toy Shape Functions and Evolution of Particle Creation Pressure
In the previous discussion, we demonstrated how the presence of particle creation mechanism at the throat provides for a possibility of a stable wormhole geometry. We utilized a inverse powerlaw variation of the particle creation pressure along the wormhole structure and derived a general shape function describing the geometry of the wormhole. Now, in this section, we are going to study the reverse process where we assume some popularly used shape functions and investigate how the particle creation pressure should vary within the wormhole. There have been several works on possibility of wormholes in GR through the use of wormhole shape functions [75, 76, 77]. Specifically we deal with three shape functions (i) $b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$, (ii) $b(r)=r_{0}^{n}r^{1-n},$ and (iii) $b(r)=r\exp[\Omega(r-r_{0})]$. The detailed analysis is carried out below:
V.1 Shape function 1: $b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$
@@
which clearly is a positive function for all values of $r$. In Fig.4, $b^{\prime}(r)$, $b(r)/r$, $b(r)-r$ and $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}$ are plotted for different values of $\epsilon$, $\xi$ and $\beta$. As required by a viable wormhole structure, the conditions mentioned in section III are clearly obeyed by the assumed shape function.
Figure 4: The viability conditions related to the shape function (37) is shown for three values of $\epsilon$ and $\beta$.
@@
Before proceeding further, let us first determine the constraints on the allowed values of $\beta$, through the analysis of squared sound speed. The sound speed is found to be
$v_{s}^{2}=\beta-\frac{8\pi}{3}$ (42)
It is found that the expression for sound speed remains unchanged, given by Eq. (27) irrespective of the shape function (and hence $\rho$, $p_{r}$ and $p_{t}$). Thus we see that the sound speed is strictly dictated by the parameter $\beta$ only. This results in the same constraints on $\beta$ given by Eq. (28). Thus, we may use the same values of $\beta$ just like the previous case, in the analysis concerned with the toy shape functions.
Now, the quantities related to the energy conditions are found to be
@@
Figure 5: The physical parameters and the energy conditions for the shape function (37) is shown for three values of $\epsilon$ and $\beta$.
From Fig. 5, it is observed that the energy density $\rho$ for this model in the presence of particle creation decreases radially, with maximum value at the throat. The evolution of the radial pressure $p_{r}$ is observed, with maximum negative values at the throat with asymptotically vanishing at far regions. This negative pressure may be thought of as a means of preventing the radial collapse of the wormhole at the throat. However, the tangential pressure remains positive. NEC is violated in terms of radial pressure at the throat but tends to obey at the asymptotic limit, whereas the DEC is obeyed at the throat as well as at the asymptotic limit. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is obeyed at the throat. Also, as $\rho\geq 0$, WEC is violated for the radial pressure case and obeyed for the tangential pressure one. Moreover, from Eq. (41) and Fig. 6, it is evident that the particle creation pressure is negative at the throat all the way to $r\to\infty$, which is required and expected for backreaction pressure responsible for sustaining a wormhole geometry. To further examine the behaviour of the effective matter resulting from the particle creation mechanism, the radial EoS parameter $\omega_{r}=p_{r}/\rho$, is also plotted that evolves with negative values, that shows exotic matter like behaviour at the throat while taking higher negative values away from the throat.
@@ V.2 Shape function 2: $b(r)=r_{0}^{n}r^{1-n}$
where $n$ is a model parameter and $r_{0}$ is the throat radius. This type of shape function has been studied in [72], in the context of $f(R)$ gravity reconstruction for given choices of wormhole shape functions. In Fig. 7 $b(r)$, $b(r)/r$, $b(r)-r$ and $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}$ are plotted for different values of $n$. It is seen that the required criteria for validity of a wormhole shape function are satisfied for different combinations of $n$.
Figure 7: The viability conditions related to the shape function (47) is shown for three values of $n$ and $\beta$.
@@
Figure 8: The physical parameters and the energy conditions for the shape function (47) is shown for three values of $n$ and $\beta$.
@@
From Fig. 8, it is seen that the energy density $\rho$ for this model in the presence of particle creation decreases radially, with maximum value at the throat. However in this case, the rate of decrease of energy density is significantly affected by the variation of $\beta$ and $n$, but the effect of $\beta$ and $n$ on the evolution of the radial pressure $p_{r}$ is noticeable but not very drastic. Similar to the previous shape functions, it takes on negative values with maximum negative values at the throat which as mentioned earlier, prevents the radial collapse of the wormhole at the throat. However, the tangential pressure remains positive and is significantly sensitive to the values of $\beta$ and $n$. The NEC is violated in terms of radial pressure at the throat but tends to obey as $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is violated at the throat and obeyed at $r\to\infty$. Also, WEC is violated in terms of the radial pressure but obeyed in terms of the tangential pressure. Moreover, from Eq. (51) and Fig. 9 (left), it is seen that the particle creation pressure is negative at the throat to $r\to\infty$. Interestingly, in this case of the shape function (56), the effective matter governed by the radial EoS parameter $\omega_{r}=p_{r}/\rho$ displays a unique behaviour contrasting to the previous cases. Although it behaves like an exotic fluid, it remains constant throughout the entire wormhole.
V.3 Shape function 3: $b(r)=r\exp[\Omega(r-r_{0})]$
@@
where $\Omega$ is the only model parameter and $r_{0}$ being the throat radius. This type of shape function has been considered widely in Ref. [73, 79, 80] in both GR and modified gravity. In Fig. 10, $b^{\prime}(r)$, $b(r)/r$, $b(r)-r$ and $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}$ are plotted for different values of $\Omega$, and shows that the validity of the throat, flaring-out and asymptotic flatness conditions are not satisfied.
Figure 10: The viability conditions related to the shape function (56) is shown for three values of $n$ and $\beta$.
However for the sake of completeness, let us calculate the physical quantities and analyze the energy conditions. The energy density corresponding to this shape function is
@@
$p_{r}=\frac{\beta(\omega+1)\left(e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}\right)}{8\pi r^{2}}-\frac{e^{-(\Omega(r-r_{0}))}}{r^{2}},$ (58)
$p_{t}=\frac{e^{\Omega(r_{0}-r)}(4\pi r\Omega-\beta(\omega+1)(r\Omega-1))}{8\pi r^{2}},$ (59)
@@
$\Pi=-\frac{\beta(\omega+1)\left(e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}\right)}{8\pi r^{2}}.$ (60)
@@
$\rho+p_{r}:=-\frac{e^{\Omega(r_{0}-r)}((\beta\omega+\beta+1)(r\Omega-1)+8\pi)}{8\pi r^{2}},$ (61)
@@
Figure 11: The physical parameters and the energy conditions for the shape function (56) is shown for three values of $\Omega$ and $\beta$.
From Fig. 11, it is seen that the energy density $\rho$ for this model in the presence of particle creation decreases radially, with maximum value at the throat. In this case, the rate of decrease of energy density is very lightly affected by the variation of $\beta$ and $\Omega$. The effect of $\beta$ and $\Omega$ on the evolution of the radial pressure $p_{r}$ is also seen to be similar but takes on negative values with maximum at the throat. With the same justification provided in the earlier cases, this negative radial pressure plays a role in preventing the radial collapse of the wormhole at the throat. However, the tangential pressure remains positive and is apparently sensitive to the values of $\beta$ and $\Omega$. We can observe NEC violation in terms of radial pressure at the throat but is obeyed at $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is satisfied at the throat as well as at asymptotic regions but DEC remains violated at the throat but obeyed at $r\to\infty$. Moreover, WEC is violated for the radial pressure case and satisfied for the tangential pressure one. Also, from Eq. (60) and Fig. 9 (right), it is seen that the particle creation pressure is negative at the throat and tends to zero at $r\to\infty$. The particle creation pressure for all the cases takes negative values at the throat which signify that although the perfect fluid matter content obeys the NEC (due to dust like EoS we considered), but the overall violation of NEC is provided by the negative particle creation pressure at the throat. Finally, the radial EoS parameter $\omega_{r}=p_{r}/\rho$, is found to evolve with negative values, that shows exotic matter like behaviour at the throat but takes on higher negative values away from the throat. Although the physical parameters seems to behave well, the overall structure of a wormhole with the shape function (56) is not allowed in the framework of particle creation, as the viability conditions are broken as discussed above.
V.4 Matching Conditions of Derived Shape Function with the Toy Shape Functions: Comparative Analysis
Let us now make a comparison of the shape function (22), derived from the physical judgements of particle creation with the toy shape functions (37), (47) and (56). If we need to estimate the explicit expression for $\alpha$ and $K$ from the shape function (22), with reference to the toy models (37), (47) and (56), we may equate the proposed general form of particle creation pressure given by Eq. (21) with the derived particle creation pressures for each toy model, and then solve for $\alpha$ and $K$. The resulting expressions for $\alpha$ and $K$ will then determine the conditions under which the toy shape functions will exactly mimic our derived shape function (22). The estimated general forms for $\alpha$ and $K$ corresponding to each toy model are tabulated in Tab. 1.
$b(r)$ $\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$ $r_{0}^{n}r^{1-n}$ $r\exp\left((r-r_{0})\Omega\right)$
$\alpha$ $\frac{\log(8\pi)-\log\left(-\frac{\beta r_{0}(\omega+1)\epsilon\coth(r_{0}\epsilon)\text{sech}^{2}(r\epsilon)}{kr^{2}}\right)}{\log(r)}$ $\frac{\log(8\pi)-\log\left(\frac{\beta(n-1)(\omega+1)r^{-n-2}r_{0}^{n}}{k}\right)}{\log(r)}$ $-\frac{\log\left(-\frac{\beta(\omega+1)e^{(r-r_{0})\Omega}(r\Omega+1)}{8\pi kr^{2}}\right)}{\log(r)}$
$K$ $-\frac{\beta r_{0}(\omega+1)\epsilon r^{\alpha-2}\text{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)}{8\pi}$ $\frac{\beta(n-1)(\omega+1)r_{0}^{n}r^{\alpha-n-2}}{8\pi}$ $-\frac{\beta(\omega+1)r^{\alpha-2}(r\Omega+1)e^{\Omega(r-r_{0})}}{8\pi}$
Table 1: Estimated general expressions for $\alpha$ and $K$ in the derived shape function (22) after matching with the toy shape functions.
VI Conclusion
@@
The concept of gravitational particle creation phenomena of open thermodynamic cosmological systems, was proposed as a potential explanation to the late-time accelerated expansion of the Universe and inflationary expansion during the early stages of the Universe. This mechanism involves the generation of repulsive gravitational effects, resulting in negative pressure, which could explain the accelerated expansion of the Universe in its later stages. Gravitational particle creation allows for speculation about violations of the NEC, which is crucial for the maintenance of a wormhole structure. This study explores the hypothesis that the geometry of a wormhole can be supported by the negative pressure created by particles. The initial section of the paper focuses on deriving the shape function of a wormhole based on a specific form of particle creation pressure. Specifically, we proposed an inverse-powerlaw form of particle creation pressure with the radial coordinate, with a source function that is linear in the Hubble parameter at the current time, denoted by $\Gamma=3\beta H_{0}$ with $\beta$ as a free parameter. Through out the paper, owing to the finiteness of the redshift function, we considered a constant redshift function for the analysis. The shape function obtained seems to adhere to the required criteria for a viable wormhole structure and violates the NEC with a positive energy density. Also the 2D and 3D embedding for the derived wormhole solution is also carried out. In the later part of the paper, in the context of particle creation mechanism we obtained the radial variation of energy density, radial and tangential pressure, and the particle creation pressure within the wormhole geometry, by assuming a few well known shape functions given by Eqs. (37), (47) and (56). The radial NEC is found to be violated at the throat in the first two case, which resonates with the behaviour of so-called exotic matter but without the need of introducing explicit form of exotic matter density like chaplygin gas [81, 69, 82, 83, 84, 85, 86, 87] or phantom matter [88, 89, 90, 91, 92, 93, 94, 95, 96]. But in the third case, the viability conditions are violated, which therefore can be ruled out as a viable shape function in the framework of particle creation. Finally, we obtained the possible general forms of $\alpha$ and $K$, the parameters those dictate the role of particle creation pressure. These general expressions represents the conditions under which the derived wormhole shape function shall mimic the assumed toy shape functions.
At this point, we may compare our study with the one carried out by Pan and Chakraborty [97] where they have studied dynamical evolving wormholes in the background of thermodynamical particle creation. They showed that evolving wormhole solutions can be obtained when the nature of the perfect fluid is phantom in nature. Contrastingly, we have found that even in the presence of a normal fluid, wormhole formation is possible.
@@
This study explores the possibility for developing a viable wormhole geometry supported by the back reaction pressure generated from the irreversible particle creation, within an open thermodynamic system. Nevertheless, our work does not address the effect of particle creation on the timelike or null geodesics within the wormhole structure or related aspects, as it is beyond the current study’s scope. This prospect will be further investigated in the future work.